OpenStudy (anonymous):

What are the possible values of x in -11x^2 + 2x = 10?

OpenStudy (anonymous):

are you looking for complex values?

OpenStudy (anonymous):

because otherwise there are none!

OpenStudy (anonymous):

\(\Large -11x^2+2x=10 \Leftrightarrow -11x^2+2x-10=0\) Now, use The Quadratic Formula : For \(\Large ax^2 + bx + c = 0\), the value of \(\Large x\) is given by: \(\Large x = \frac{ -b ± \sqrt{b^2 - 4ac} }{ 2a}\)

OpenStudy (anonymous):

Refer to the attached plot for x ranging from -1000 through 1000.